# A Weibull Mixture Cure Frailty Model for High-dimensional Covariates

### v1

@@ 5.2 Functional enrichment analysis
In Scenario $1$, only one biological process (BP) for the GO-EA is significant, the “xenobiotic metabolic process”, and the association between this BP and similar breast cancer evolution processes is examined by Lee et al. [48]. Another BP identified by GO-EA is “GABAergic neuron differentiation”, with a p-value of $0.057$. For KEGG-EA, the only significantly enriched term is “metabolism of xenobiotics by cytochrome P450”. The association between this enzyme expression and cancer risk, progression, metastasis and prognosis has been widely reported in basic and clinical studies (see the review study of Luo et al. [62] for more details). Also the role of CYP1A1 and CYP2A13 genes in breast and lung cancers has been investigated extensively (see Sneha et al. [82] and references therein for more details). The gene network plots related to the significant GO and KEGG pathway terms are presented in Figures 6(a) and 6(b). The EA results for other methods, namely penMCFM(GMIFS) and MCM(GMIFS), are respectively shown in Figures S2222-S3333 in the Supplementary Materials. While the number of enriched terms in both GO- and KEGG-EA is limited for all the methods, some of the identified terms exhibit relevance to breast cancer.
In Scenario $2$, we detect totally $28$ enriched BP terms via GO-EA, including some important BPs associated with breast cancer, such as “Notch signaling pathway” [102], “canonical Wnt signaling pathway” [97, 1], and “BMP signaling pathway” [29]. The barplot for some of the enriched terms in GO-EA, and the gene network plot related to the top-significant GO terms, are presented in Figures 7(a) and 7(b) respectively. Concerning results of KEGG-EA in this Scenario, Figure 8(a) shows that the “Wnt signaling” and “breast cancer pathways” are two of the most significant KEGG pathways, as confirmed by the growing number of studies in the literature demonstrating that Wnt signaling involves the proliferation, metastasis, immune microenvironment regulation, stemless maintenance, therapeutic resistance, and phenotype shaping of breast cancer (see Xu et al. [97] and Abreu de Oliveira et al. [1], and references therein, for more details). The gene network plot related to the significant KEGG-EA pathway terms is presented in Figure 8(b). Moreover, the enriched KEGG pathway “hsa05224: Breast cancer” is reported in Figure S5555 in the Supplementary Materials to more clearly show the associated biological meaning. We observe a greater number of enriched terms in the EA for all methods in Scenario $2$ compared to Scenario $1$. Notably, the “breast cancer” pathway is identified as one of the most enriched terms in at least one part of the model structure for all methods.
@@ 5.3 Prognostic risk score construction and validation of identified genes
We use the validation dataset, which includes $189$ observations with their RNA-seq features and survival time, to validate the selected genes with the proposed PRS. We assign the subjects in the validation dataset to a high and low risk group using the median of the obtained PRS values. This procedure is applied when using the selected genes obtained with penMCFM(EM), penMCFM(GMIFS), MCM(GMIFS) and penCox.1se in Scenario $2$. The KM curves of these subgroups are presented in Figure 10 for penMCFM(EM) and MCM(GMIFS). Since the $p$-values of the log-rank tests for penMCFM(GMIFS) and penCox.1se are not significant, these KM plots are only shown in Figures S14141414(a)-(b) in the Supplementary Materials. From Figure 10, it can be observed that there are significant differences between the survival distributions corresponding to the high and lower risk groups as obtained by these two methods. When considering the obtained log-rank test p-values, $0.018$ and $0.027$ for penMCFM(EM) and MCM(GMIFS), respectively, and examining the survival curves in the figures, the subgroups defined by penMCFM(EM) appear to be more distinguishable than those defined by MCM(GMIFS), thus showing a better prognostic power of the former proposed method.
Moreover, the heatmaps of the expression values of the selected genes, based on the penMCFM(EM) and MCM(GMIFS) methods, in the validation dataset are also presented in Figure S15151515 in the Supplementary Materials. In these heatmaps, only the genes which are selected at least $2(3)$ out of $20$ repeats are shown, for visualization purposes. It can be seen from inspection of these figures that the gene clusters based on the penMCFM(EM) method are more visible than those based on the MCM(GMIFS) method.
@@ 6 Discussion and Conclusion
In this study, we have introduced a novel method that adapts the mixture cure frailty model to account for high-dimensional covariates in both the incidence and the latency components of the model. The method extends the classical MCFM to higher dimensions using a multi-step adaptive elastic net penalty term inside the EM algorithm, therefore the name penMCFM. The proposed penMCFM allows performing variable selection in the MCFM model in high-dimensional settings where the number of covariates is significantly larger than the sample size. A comparative analysis with alternative survival models, including the penalized mixture cure model and penalized Cox regression, was also carried out. The numerical results presented in the simulation studies indicate that penMCFM outperforms the alternative approaches under study with respect to various metrics of prediction and estimation performances. We also applied the proposed model along with alternative approaches to analyze RNAseq data from BRCA samples from TCGA. Finally, we conducted functional enrichment analyses of the identified biomarkers to show the disease mechanisms associated with BRCA, and then validated these biomarkers using an independent validation dataset to demonstrate their effectiveness and usefulness.
@@
Some potential directions for future research include exploring a semi-parametric approach as an alternative to the current use of parametric distributions in the latency part. Additionally, investigating a Bayesian approach for variable selection could be a fruitful direction for further investigation.
@@ Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 801133. The results published in the real data study are based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga.
@@
For the incidence part, we detect only one significant enriched KEGG pathway, namely the “hsa04726: Serotonergic synapse”. For the latency part, we have not identified any enriched KEGG-EA pathway terms, the barplot of significantly enriched GO terms is given in Figure S3.
Figure S3: Results of the analysis of the TCGA-BRCA data. Barplot of significantly enriched GO terms for the latency part of the model based on MCM(GMIFS).
@@
Figure S15: Results of the analysis of the TCGA-BRCA data. Heatmap of the expression values of the selected genes using the average results of $20$ repeats in Scenario $2$ for the validation dataset: (a) penMCFM(EM),genes that are selected in at least $2$ over $20$ repeats, and (b) MCM(GMIFS), genes that are selected in at least $3$ over $20$ repeats
@@
Biomarker
Description
Reported literature
KEGG pathway
IGLL5
@@
The IGLL5 gene can serve not only as a prognostic biomarker for breast cancer patients but also as a potential source of therapeutic strategies for preventing breast cancer recurrence. Its expression is also correlated with tumor‐infiltrating immune cells.
[5, 50, 59, 93]
GABRQ
@@
It has been demonstrated that the mRNA expression of GABRQ could serve as a prognostic marker for clear cell renal cell carcinoma and colon adenocarcinoma. It has been identified as one of the upregulated genes in triple-negative breast cancer (TNBC).
[49, 99, 54]
L1CAM
@@
L1CAM is expressed in numerous human cancers and is frequently associated with poor prognosis. It is suggested that L1CAM can be utilized for breast cancer diagnosis, indicating a potential correlation between L1CAM expression and the overall adverse prognosis of TNBC.
[26, 2, 6]
ADGRG7
@@
The role of this gene in head and neck squamous cell carcinoma, uterine corpus endometrial carcinoma, and breast cancers are investigated. It is observed that the gene is directly involved in breast tumor metastasis to bone tissues.
[67, 51, 81]
VAX1
@@
It has been identified as one of the $10$ upregulated genes in TNBC. Its association with lung squamous cell carcinoma has also been observed, suggesting its potential as a prognostic biomarker for evaluating risk assessment.
[37, 56]
CSAG3
@@
CSAG3, also known as TRAG-3 (Taxol Resistance Associated Gene-3), is reported to be upregulated in numerous tumors, including gastric cancer, urothelial carcinoma of the bladder, ovarian carcinoma, and melanoma.
[105]
ENHO
@@
Adropin is encoded by the energy homeostasis-associated (ENHO) gene. Its association with the development of colorectal cancer has also been investigated. GPR19 is activated by adropin is studied for the breast tumor cells.
[77, 45]
OR56A3
@@
The Olfactory Receptor (OR) family has gained attention as a potential biomarker for cancer. The significance of transcript abundance in certain OR genes is examined in an invasive breast carcinoma population. However, a substantial portion of the roles of OR genes in breast cancer remains understudied.
[90, 65]
C20orf85
@@
It has been implicated in various cancer types, including lung cancer, lower-grade glioma of the brain, ovarian cancer, and breast cancer. Further investigations are needed to explore its implications.
[41, 79, 11, 36]
GREB1L
@@
It has been observed that the GREB1L gene is implicated in the development of breast cancer, and it is proposed as a potential molecular marker for predicting the prognosis of breast cancer. It has been also demonstrated to exhibit a high correlation with both estrogen receptor and androgen receptor expression in breast/prostate cancer cell lines and primary tumors.
[12, 27]
BMPR1A
@@
The Bone Morphogenetic Protein Receptor (BMPR) genes associated with various cancer types is explored. The role of the BMPR1A gene in breast cancer growth and metastasis is also investigated.
[70, 75, 40]
TSPEAR
@@
The TSPEAR gene has not been reported in any cancer study, except for one preprint study related to colorectal cancer. Further investigations are required to delve into its implications.
[58]
KRT77
@@
The differential expression of KRT genes is investigated across various cancers. Its potential role as a biomarker in head and neck squamous cell carcinoma is also explored.
[25, 85]
@@
Biomarker
Description
Reported literature
KEGG pathway
GNG4
It has been reported as a potential biomarker in various cancer types, including bladder, colorectal, gastric, and breast cancer.
[106, 64, 28, 6]
SERPINB7
The SERPINB family genes are differentially expressed in the tumor tissues. The SERPINB7 gene has been studied in various cancer types, including breast, cervical, and lung cancers.
[20, 91, 69]
GABRG2
GABA receptor genes (GABR) constitute a group of genes associated with developmental and epileptic encephalopathies. They have also been identified to be associated with recurrent breast, colon, and laryngeal cancer samples.
[43, 86, 99, 68]
ZNF385B
The potential effects of ZNF385B expression in breast cancer have been determined in recent studies, suggesting its utility as a potential diagnostic and prognostic biomarker for breast cancer. It is also observed to be correlated with the overall survival of ovarian cancer patients.
[30, 100, 110]
CT83
CT83 is highly expressed in gastric, triple-negative breast, lung, and hepatocellular cancers. It is observed to be significantly associated with the overall survival of TNBC patients.
[109, 19, 55]

### v2

@@ 5.2 Functional enrichment analysis
In Scenario $1$, only one biological process (BP) for the GO-EA is significant, the “xenobiotic metabolic process”, and the association between this BP and similar breast cancer evolution processes is examined by Lee et al. [60]. Another BP identified by GO-EA is “GABAergic neuron differentiation”, with a p-value of $0.057$. For KEGG-EA, the only significantly enriched term is “metabolism of xenobiotics by cytochrome P450”. The association between this enzyme expression and cancer risk, progression, metastasis and prognosis has been widely reported in basic and clinical studies (see the review study of Luo et al. [61] for more details). Also the role of CYP1A1 and CYP2A13 genes in breast and lung cancers has been investigated extensively (see Sneha et al. [62] and references therein for more details). The gene network plots related to the significant GO and KEGG pathway terms are presented in Figures 7(a) and 7(b). The EA results for other methods, namely penMCFM(GMIFS) and MCM(GMIFS), are respectively shown in Figures S2222-S3333 in the Supplementary Materials. While the number of enriched terms in both GO- and KEGG-EA is limited for all the methods, some of the identified terms exhibit relevance to breast cancer.
In Scenario $2$, we detect totally $28$ enriched BP terms via GO-EA, including some important BPs associated with breast cancer, such as “Notch signaling pathway” [63], “canonical Wnt signaling pathway” [64, 65], and “BMP signaling pathway” [66]. The barplot for some of the enriched terms in GO-EA, and the gene network plot related to the top-significant GO terms, are presented in Figures 8(a) and 8(b) respectively. Concerning results of KEGG-EA in this Scenario, Figure 9(a) shows that the “Wnt signaling” and “breast cancer pathways” are two of the most significant KEGG pathways, as confirmed by the growing number of studies in the literature demonstrating that Wnt signaling involves the proliferation, metastasis, immune microenvironment regulation, stemless maintenance, therapeutic resistance, and phenotype shaping of breast cancer (see Xu et al. [64] and Abreu de Oliveira et al. [65], and references therein, for more details). The gene network plot related to the significant KEGG-EA pathway terms is presented in Figure 9(b). Moreover, the enriched KEGG pathway “hsa05224: Breast cancer” is reported in Figure S5555 in the Supplementary Materials to more clearly show the associated biological meaning. We observe a greater number of enriched terms in the EA for all methods in Scenario $2$ compared to Scenario $1$. Notably, the “breast cancer” pathway is identified as one of the most enriched terms in at least one part of the model structure for all methods.
@@ 5.3 Prognostic risk score construction and validation of identified genes
We use the validation dataset, which includes $189$ observations with their RNA-seq features and survival time, to validate the selected genes with the proposed PRS. We assign the subjects in the validation dataset to a high and low risk group using the median of the obtained PRS values. This procedure is applied when using the selected genes obtained with penMCFM(EM), penMCFM(GMIFS), MCM(GMIFS) and penCox.1se in Scenario $2$. The KM curves of these subgroups are presented in Figure 10 for penMCFM(EM) and MCM(GMIFS): in the figure it can be observed that the log-rank test shows significant differences between the survival distributions corresponding to the high- and low-risk groups as obtained by the PRS calculations carried out on the outputs of both penMCFM(EM) and MCM(GMIFS) methods (with p-values $0.018$ and $0.035$, respectively). Since the $p$-values of the log-rank tests for penMCFM(GMIFS) and penCox.1se are not significant, these KM plots are only shown in Figures S14141414(a)-(b) in the Supplementary Materials.
Moreover, the heatmaps of the expression values of the selected genes, based on the penMCFM(EM) and MCM(GMIFS) methods, in the validation dataset are also presented in Figure S15151515 in the Supplementary Materials. In these heatmaps, only the genes which are selected in at least $2$ out of $20$ re-run of the methods are shown, for visualization purposes. From inspection of these figures it can be observed that the penMCFM(EM) method gives much more visible clusters of genes showing similar expression behaviours than those based on the MCM(GMIFS) method.
@@ 6 Discussion and Conclusion
In this study, we have introduced a novel method that adapts the mixture cure frailty model to account for high-dimensional covariates in both the incidence and the latency components of the model. The method extends the classical MCFM to higher dimensions using a multi-step adaptive elastic net penalty term inside the EM algorithm, therefore the name penMCFM. The proposed penMCFM allows performing variable selection in the MCFM model in high-dimensional settings where the number of covariates is significantly larger than the sample size. A comparative analysis with alternative survival models, including the penalized mixture cure model and penalized Cox regression, was also carried out. We also applied the proposed model along with alternative approaches to analyze RNAseq data from BRCA samples from TCGA. Finally, we conducted functional enrichment analyses of the identified biomarkers to show the disease mechanisms associated with BRCA, and then validated these biomarkers using an independent validation dataset to demonstrate their effectiveness and usefulness.
Moreover, we conducted a comprehensive investigation comparing our primary methodology, penMCFM(EM), with alternative methods using various metrics including C-statistic, AUC, and IBS, alongside functional enrichment analysis and the construction of prognostic risk scores, on various real data scenarios. Our analysis revealed that penMCFM(EM) and MCM(GMIFS) methods demonstrated comparable performances under certain circumstances, while penMCFM(EM) exhibited superior performance in others. Notably, despite similarities in performance observed in certain scenarios, penMCFM(EM) consistently selected fewer genes than MCM(GMIFS), its main competitor. Moreover, through rigorous simulation studies, penMCFM(EM) consistently outperformed alternative methods across different evaluation metrics. Consequently, we conclude that penMCFM(EM) represents a compelling alternative, offering advantages across diverse data generating scenarios. Additionally, our future research agenda entails applying this method to other cancer datasets, and conducting further comparative analyses against relevant methods.
@@
Some potential directions for future research include exploring a semi-parametric approach as an alternative to the current use of parametric distributions in the latency part. Additionally, investigating a Bayesian approach for variable selection could be a fruitful direction for further investigation. Finally, to the best of our knowledge, the literature on cure frailty models for recurrent events in high-dimensional settings has not flourished to the same extent as that based on survival data. This could be another potentially fruitful direction of future research.
@@ Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 801133. Computer simulations were performed in HPC solutions provided by the Oslo Centre for Epidemiology and Biostatistics, at the University of Oslo. The results published in the real data study are based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga.
@@
For the incidence part, we detect only one significant enriched GO term, namely the “GO:0001533 cornified envelope”, and no significant enrichment terms are identified in KEGG-EA. For the latency part, the significantly enriched GO-EA and KEGG-EA pathway terms are given in Figure S3.
Figure S3: Results of the analysis of the TCGA-BRCA data. GO-EA and KEGG-EA for the latency part of the model based on MCM(GMIFS): (a) Barplot of significantly enriched GO terms (b) Network plot of enriched KEGG pathway terms and related selected genes
@@
Figure S15: Results of the analysis of the TCGA-BRCA data. Heatmap of the expression values of the selected genes that are selected in at least $2$ over $20$ repeats in Scenario $2$ for the validation dataset: (a) penMCFM(EM) and (b) MCM(GMIFS)
@@
Biomarker Description Reported literature KEGG pathway
@@
IGLL5 The IGLL5 gene can serve not only as a prognostic biomarker for breast cancer patients but also as a potential source of therapeutic strategies for preventing breast cancer recurrence. Its expression is also correlated with tumor‐infiltrating immune cells. [67, 68, 69, 70]
@@
GABRQ It has been demonstrated that the mRNA expression of GABRQ could serve as a prognostic marker for clear cell renal cell carcinoma and colon adenocarcinoma. It has been identified as one of the upregulated genes in triple-negative breast cancer (TNBC). [71, 72, 73]
@@
L1CAM L1CAM is expressed in numerous human cancers and is frequently associated with poor prognosis. It is suggested that L1CAM can be utilized for breast cancer diagnosis, indicating a potential correlation between L1CAM expression and the overall adverse prognosis of TNBC. [74, 75, 76]
@@
ADGRG7 The role of this gene in head and neck squamous cell carcinoma, uterine corpus endometrial carcinoma, and breast cancers are investigated. It is observed that the gene is directly involved in breast tumor metastasis to bone tissues. [77, 78, 79]
@@
VAX1 It has been identified as one of the $10$ upregulated genes in TNBC. Its association with lung squamous cell carcinoma has also been observed, suggesting its potential as a prognostic biomarker for evaluating risk assessment. [80, 81]
@@
CSAG3 CSAG3, also known as TRAG-3 (Taxol Resistance Associated Gene-3), is reported to be upregulated in numerous tumors, including gastric cancer, urothelial carcinoma of the bladder, ovarian carcinoma, and melanoma. [82]
@@
ENHO Adropin is encoded by the energy homeostasis-associated (ENHO) gene. Its association with the development of colorectal cancer has also been investigated. GPR19 is activated by adropin is studied for the breast tumor cells. [83, 84]
@@
OR56A3 The Olfactory Receptor (OR) family has gained attention as a potential biomarker for cancer. The significance of transcript abundance in certain OR genes is examined in an invasive breast carcinoma population. However, a substantial portion of the roles of OR genes in breast cancer remains understudied. [85, 86]
@@
C20orf85 It has been implicated in various cancer types, including lung cancer, lower-grade glioma of the brain, ovarian cancer, and breast cancer. Further investigations are needed to explore its implications. [87, 88, 89, 90]
@@
GREB1L It has been observed that the GREB1L gene is implicated in the development of breast cancer, and it is proposed as a potential molecular marker for predicting the prognosis of breast cancer. It has been also demonstrated to exhibit a high correlation with both estrogen receptor and androgen receptor expression in breast/prostate cancer cell lines and primary tumors. [91, 92]
@@
BMPR1A The Bone Morphogenetic Protein Receptor (BMPR) genes associated with various cancer types is explored. The role of the BMPR1A gene in breast cancer growth and metastasis is also investigated. [93, 94, 95]
@@
TSPEAR The TSPEAR gene has not been reported in any cancer study, except for one preprint study related to colorectal cancer. Further investigations are required to delve into its implications. [96]
@@
KRT77 The differential expression of KRT genes is investigated across various cancers. Its potential role as a biomarker in head and neck squamous cell carcinoma is also explored. [97, 98]
@@
Biomarker Description Reported literature KEGG pathway
GNG4 It has been reported as a potential biomarker in various cancer types, including bladder, colorectal, gastric, and breast cancer. [99, 100, 101, 76]
SERPINB7 The SERPINB family genes are differentially expressed in the tumor tissues. The SERPINB7 gene has been studied in various cancer types, including breast, cervical, and lung cancers. [102, 103, 104]
GABRG2 GABA receptor genes (GABR) constitute a group of genes associated with developmental and epileptic encephalopathies. They have also been identified to be associated with recurrent breast, colon, and laryngeal cancer samples. [105, 106, 72, 107]
ZNF385B The potential effects of ZNF385B expression in breast cancer have been determined in recent studies, suggesting its utility as a potential diagnostic and prognostic biomarker for breast cancer. It is also observed to be correlated with the overall survival of ovarian cancer patients. [108, 109, 110]
CT83 CT83 is highly expressed in gastric, triple-negative breast, lung, and hepatocellular cancers. It is observed to be significantly associated with the overall survival of TNBC patients. [111, 112, 113]